find the slope of the line (x/4)-(y/5)=1
Answer by mananth(16949)

x/4 - y/5 =1
LCD = 20
..
5x-4y /20 =1
5x-4y= 20
add -5x to both sides
5x-5x-4y=20-5x
-4y=-5x+20
divide by -4
-4y/-4=-5x/-4 -20/-4
y=5x/4 +5
compare with y = mx +b
m= 5/4 the slope